My first title

Eirini

March 24, 2022

1 My first section

My first paragraph

1.1 My first subsection

”LaTeX is a \underline{document} preparation system and document markup language. LaTeX uses the TeX typesystem program for formating its putput, and is

itself written in TeX macro language”.

LaTeX is not the name of a particular editing program, but refers to the encoding or tagging conventions that are used in LaTeX documents.